Status Report

NASA Space Station On-Orbit Status 7 Aug 2003

By SpaceRef Editor
August 7, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 7 Aug 2003

All ISS systems continue to function nominally, except as noted previously or below.  Week 14 of Expedition 7.  As of this morning, the station has circled Earth 26,900 times since first launch (FGB/Zarya), covering a distance of 1.143 billion km (714.4 million miles).  That’s 7.6 times the distance from us to the Sun.

Main activities of the crew today were checkout, resizing, and fitcheck of one of the three available EMUs (extravehicular mobility units) for Science Officer Ed Lu.  Using the extra-large spacesuit #3005, the plan was to first verify proper cooling flow, followed by resizing and finally fit verification.  The latter two objectives were successfully accomplished, but cooling at first was intermittent before it finally stabilized OK.  This needs to be investigated, and more troubleshooting time will be scheduled next week.  [On 5/28, during the two-man EVA (extravehicular activity) don/doff dryrun, Ed was unable to establish cooling water flow for his EMU #3013.  Today, suit #3005 was resized by swapping arms and LTAs (lower torso assemblies) between it and #3013.  Due to concern about possible EMU cooling water contamination in the LCVGs (liquid cooling ventilation garments), Ed used the LCVG worn by Malenchenko during the don/doff dryrun, which had been filled on orbit with water containing iodine (the other two garments, filled on the ground with de-ionized water, being more prone to developing contamination growth).  For the same reason, SCU-2 (servicing & cooling umbilical #2) was used today instead of SCU-1.]

Before the EMU activities could be started, MCC-Moscow disabled the 400 MHz (megahertz) transmitter (PRD) of the Russian/German Global Timing System (GTS).  Later, after EMU ops were completed, GTS was switched on again.  [GTS electronics are installed in the Service Module (SM), with two transmitters broadcasting a time signal at 400.1 MHz (1 watt) and 1428 MHz (0.5 watt) straight downward (nadir) for worldwide clock synchronization.  Because the 400 MHz signal interferes with the WVS (Wireless Video System) and very likely also the SSCS (Space-to-Space Communication System), both used for spacewalks, the GTS 400 MHz transmitter must be inhibited during EVAs, including EMU checkouts, from the ISS Joint Airlock (A/L).]

After the EMU activities, Ed Lu reconfigured the flexible ventilation duct in the A/L to IMV (intermodular ventilation) air return, and MCC-H deactivated the A/L’s air conditioner CCAA (common cabin air assembly), which ran since yesterday.

Analysis of pictures taken by Ed during his A/L closeout panel survey on 5/26 for broken fasteners, current panel configuration and any configuration issues indicates some concern with “kick loads” on two of the panels.  The crew was asked to attempt to tighten some fasteners on those two panels before beginning today’s EMU activities.  [Kick forces are sometimes inadvertently applied when donning/doffing space gear while floating in the narrow confines of the A/L.]

FE/SO Lu finished the periodic ELPS (emergency lighting power sources) inspection, begun in the Node on 8/4 and focusing today on A/L “Quest” and Lab “Destiny”.

Before beginning his scheduled EPO (Educational Payload Operations) activity session at 12:15pm EDT, Dr. Lu prepared and set up the video camcorder for taping the demonstration.  [The activity demonstrated the use of EVA and IVA (intravehicular activity) tools in space, including IVA drill driver, EVA pistol grip tool (PGT), high-torque locker tool and torque wrench, clamp and bracket kit, leak repair kit, body restraint tether use, maintenance work area, tool restraints, etc.  The resulting video will be used to supplement NASA educational materials, as well as be featured at NASA education websites.]

CDR Yuri Malenchenko continued his current (seventh) session of the Russian Uragan (“hurricane”) earth imaging program (GFI-8), using the Kodak DCS760 digital still camera with 800-mm lens where cloud cover did not exceed 30%.  [Targets included the Kura river valley along the route Borzhomi-Derbi, Kolka glacier and other glaciers on Kazbek volcano slopes, the cities of Vladikavkaz and Grozny (burning wells), Caspian Sea coastal zones, Ekibastuz quarries, and the Katun river valley.  The images, in TIF format, were again stored on the removable hard disk drive (HDD) #4 of the French EGE-2 laptop.]

Yuri also completed another photography session for the Diatomeya ocean observations program, using the DVCAM 150 digital video camera and the Nikon-F5 electronic still camera with 80-mm lens.  [Today’s observations again presented an opportunity to investigate bioproductive waters in the Northern Atlantic and its surrounding seas (Caribbean, Mediterranean and Black Sea) which are being actively researched by ships.  Specific targets included the Yucatan Channel and Florida Channel mouth, US and Canadian East Coast, Cuba shelf, Bahamas sand-banks, Sargasso Sea, etc.]

Ed completed his thirteenth weekly filling-out of the FFQ (food frequency questionnaire), which keeps a log of his nutritional intake over time on special medical equipment computer (MEC) software.

The CDR performed the regular daily routine tasks of SOZh life support system maintenance and inventory management system [IMS] delta file preparation.  Yuri also conducted the weekly inspection of the BRPK-1 air/water condensate separator.

Early in the morning, Dr. Lu collected and stowed the two passive FMK (formaldehyde monitoring kit) sampling badges which he had deployed in the Lab and SM two days ago (8/5).

Both crewmembers worked out in their daily 2.5-h program of physical exercise, on TVIS treadmill and RED expander.

There were no MSG/CSLM-2 (Microgravity Science Glovebox/Coarsening in Solid/Liquid Mixtures-2) crew activities today.  [The samples in SPU-6 (sample processing unit #6) are continuing to “soak”, for the duration of today.  During Soaking mode, the electronic control unit (ECU) cycles the internal heaters in order to maintain the samples at 185 °C.]

PCU-2 (Plasma Contactor Unit #2), switched to standby yesterday, was activated again overnight for another 24-hr. run, discharging Xenon gas to continue characterization of the plasma environment. characterization.  [The test is part of a continuing effort to establish a relationship between the charging hazard, ambient ionosphere conditions, and the current flight configuration of LVLH XVV.]

MCC-M is continuing the checkout tests of the newly installed Russian ASN-1 satellite navigation antenna system which has encountered technical problems.  Today’s tests, during Daily Orbits 2 and 3, were conducted with the ASN-1 in “GPS-only” mode.

At 10:00am, MCC-H performed the second part of the transition of the S1 and P1 truss segment MDMs to software version R2, today finishing off with S1-2 and P1-2 computers.  [The new software adds TCS (thermal control system) capability required for Mission 12A.1 and thus has no significant impacts to the crew or their procedures for the duration of Increment 7.  While some C&W (caution & warning) annunciations are expected during this process, the ground will do everything necessary to keep them suppressed.]

Starting 8/5 and running through 8/10, MCC-H and POC (Payload Operations Center) are continuing a Ku-band 150 MBPS (megabytes per second) evaluation test, during which the IP Phone and OCA transfer capabilities are temporarily out of service.  [Currently, MCC-H and POC use the 50 Mbps capability of Ku-band.  The purpose of this test is to demonstrate that the ISS Ku-band 150 Mbps and 75 Mbps can be effectively utilized for future day-to-day operations.]

MCC-M/TsUP will not have command capability from today through 8/9 (Saturday) due to maintenance at the Shelkovo ground station and unavailability of the Molniya satellite.  Voice capability will not be affected during this period.  [RGS Shelkovo telemetry equipment will be operational a day earlier, on 8/9, but command equipment will not be ready until 8/10 (Sunday).]

Also on 8/9, there will be a 3-hr. outage of Ku-band downlink capability at MCC-H, to support installation of new power (high-rate data switch unavailable), starting at 11:30am EDT.

Last Saturday (8/2), TsUP performed the periodic efficiency test on the RS (Russian segment) solar arrays.  [There were no issues, and propellant use for the test from maneuver begin to return to CMG momentum management amounted to 35 kg (77 lbs).]

At regular intervals, the station’s cabin atmosphere is refreshed with gaseous oxygen (O2) from Progress 10P O2 tanks as long as its supply lasts (originally 50 kg/110 lbs), whenever ppO2 (O2 partial pressure) decreases to 160 mmHg.  Afterwards, O2 repress will also be performed from Progress 11P.  [During the Progress-supplied repress period, the Elektron oxygen generator is off, to conserve its service life, and the repress valve cycling, by the crew, relies on the U.S. MCA (major constituents analyzer) or CSA-CP (compound specific analyzer-combustion products) instruments to measure the 10mmHg pressure difference.]

According to current TsUP planning, Progress 10P will undock in three weeks, on 8/28, followed by 11P undocking on 9/4.  These dates are still unofficial.

During the night of 8/4, as the crew slept and the solar Beta angle decreased to below 10 deg., the ISS maneuvered from XPOP (x-axis perpendicular to orbit plane) back to LVLH XVV (local vertical, local horizontal/x-axis in velocity vector).

Today’s CEO (crew earth observation) targets, no longer limited in the current LVLH attitude and including the targets of the Lewis & Clark 200-year memorial locations, were Great Wall, China (known wall remnants on the west side [‘far side of the easy-to-see river from the crew’s location] of the Huang Ho, where a color change in the grassland on either side of the wall has been observed in past handheld imagery.  Then, ~1.5 min later, a nadir view where the wall crosses the Ordos desert [in the great bend of the Huang Ho]), Cape Town, South Africa (nadir pass.  The crew’s prior interesting views have been a touch too oblique for CEO science purpose.  The 180 mm lens is a requirement for the CEO analysts to remove this target from the list), Tashkent, Uzbekistan (looking right at the foot of the mountains), Eastern Mediterranean Aerosols (visibilities somewhat reduced to 4-5 miles in local cities.  Pointing right for obliques.  This was a good opportunity to start to calibrate handheld imagery in comparison wit local visibility measurements and the TOMS satellite measurements), Alexandria, Egypt (nadir pass), Lisbon, Portugal (pointing a touch right on the major Duero River estuary), Berlin, Germany (pointing a touch left.  Due to high pressure in northern Europe, good viewing from the English Channel to Belarus), London, England (heat wave in Europe [low 90s in southern England!]), and Missouri River sites (LEWIS & CLARK SITES:  Track crossed the middle of the site  pointing left and right).
CEO images can be viewed at the websites

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 12:05pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is operating.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.

Propulsion System (PS):

  • Total propellant load available: 3646 kg (8038 lb) as of 7/31  [SM(774) + FGB(2872) + Progress M(0) +Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (blind, dual-angle mode with drag reduction, non solar-tracking).
  • SM batteries:  Battery #5 is off line(Capacity Restoration mode, ROM); all other batteries (7) are in “Partial Charge” mode.  Batteries #1and #3 showing degraded capacity. 
  • FGB batteries:  Battery #1 is off line (Capacity Restoration mode, ROM); battery #5 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Discharge mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-1 is operating; INT-2 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is Off; PL-1 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH -YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, -y-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.
  • Solar Beta angle:  3.3 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:53am EDT [= epoch]):

  • Mean altitude — 384.8 km
  • Apogee — 390.5 km
  • Perigee — 379.1 km
  • Period — 92.25 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0008414
  • Orbits per 24-hr. day — 15.61
  • Solar Beta Angle — 3.3 deg (magnitude decreasing)
  • Mean altitude loss in last 24 hours — 100 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 26900
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.